Price is £30. Quantity is 200. Price increases to £33, Quantity falls to 100.
my answer was 5 therefore it is elastic demand.
But in my test I got the question on price elasticity of demand wrong so I want to double check this is right before I send it to my teacher to make up for the test
1
Expert's answer
2020-10-16T10:41:23-0400

solution

price elasticity of demand=(%change in quantity )/ (%change in price)


% change in quantity"(\Delta Q)=\frac{(200-100)}{200}\times100=50" %


%change in price"(\Delta P)=\frac{(33-30)}{30}\times100=10" %


so


so price elasticity of demand "=" "\frac{50}{10}=5"


so price elasticity of demand is 5 .